# Re-Examine Distantly Supervised NER: A New Benchmark and a Simple Approach

### v1

@@ Abstract
This paper delves into Named Entity Recognition (NER) under the framework of Distant Supervision (DS-NER), where the main challenge lies in the compromised quality of labels due to inherent errors such as false positives, false negatives, and positive type errors. We critically assess the efficacy of current DS-NER methodologies using a real-world benchmark dataset named QTL, revealing that their performance often does not meet expectations. To tackle the prevalent issue of label noise, we introduce a simple yet effective approach, Curriculum-based Positive-Unlabeled Learning (CuPUL), which strategically starts on “easy” and cleaner samples during the training process to enhance model resilience to noisy samples. Our empirical results highlight the capability of CuPUL to significantly reduce the impact of noisy labels and outperform existing methods.
@@ A New Benchmark and a Simple Approach
Yuepei Li                        Kang Zhou                        Qiao Qiao Iowa State University {liyp0095, kangzhou, qqiao1, qingwang, qli}@iastate.edu                        Qing Wang                        Qi Li
@@ 1 Introduction
Distantly-Supervised Named Entity Recognition (DS-NER) is a task to leverage existing knowledge bases (KBs) or dictionaries to provide annotations for named entity recognition tasks. This approach significantly reduces the need for labor-intensive manual annotations, but it faces challenges due to issues in automated annotations, such as false positives, false negatives, and positive type errors. To address the annotation errors, various methods are proposed. Some studies focus on false negative issues Shang et al. (2018); Peng et al. (2019); Zhou et al. (2022). Others propose to tackle general noisy annotations through noise removal processes Meng et al. (2021); Liang et al. (2020); Hedderich and Klakow (2018); Zhang et al. (2021a); Liu et al. (2021).
DS-NER approaches have been successfully applied to existing NER benchmark datasets, achieving competitive performances with fully supervised methods. However, in our usage in a real-life DS-NER application, we observe that the existing DS-NER approaches fail to obtain satisfactory performances. In this paper, we present a new benchmark dataset named QTL and re-examine the existing DS-NER methods in a real-life DS-NER task. We further present a simple yet effective approach following the philosophy of curriculum learning.
Previous DS-NER approaches primarily rely on noise removal to address annotation issues especially false positives and positive type errors. However, these methods face several challenges when applied in real-life DS-NER tasks. First, a moderately well-trained NER model is necessary to detect noise. However, it is hard to determine proper hyper-parameters when an NER model is moderately well-trained without sufficient validation data. Stopping the training too early, the model cannot produce an accurate enough model for noise detection. Stopping the training too late, the model will learn the noise and degrade the performance. Second, the moderately well-trained NER model is trained on noisy labels initially, so the noise detection methods may have unknown biases and cause irreparable damage.
@@
Instead of removing noisy labels, we propose a simple yet robust approach using token-level Curriculum-based Positive-Unlabeled Learning (CuPUL) to tackle the challenge of noisy labels in DS-NER tasks. The motivation of curriculum learning is that deep learning models are non-convex and trained using batches of samples, so the order of training data can significantly impact the model performance. Curriculum Learning rearranges the batches of training samples such that the model learns from easy to hard and learns from easy samples more times. With the new arrangement, the models tend to converge to a better local optimum. We further design a token-level curricula arrangement to address the token-level noise for DS-NER tasks. We observe that “easy samples” are usually cleaner. Consequentially, learning from easy samples first can avoid label noise initially and make the model more robust. We adopt the Positive-Unlabeled (PU) learning paradigm to address the false negative issues.
@@
  • We present a real-life DS-NER application, QTL, and test the performance of the existing state-of-the-art methods. We observe that many methods do not follow the practical DS-NER setting and have unsatisfactory performance.
@@
  • We propose a simple method CuPUL to address the noise issue in DS-NER. We empirically demonstrate that CuPUL can significantly outperform the state-of-the-art DS-NER method on the QTL dataset.
  • CuPUL also achieves good performance on existing benchmarks.
2 Related Work
2.1 Benchmark Datasets
@@
To reduce the cost of human-annotated training data for NER tasks, DS-NER uses professional dictionaries or knowledgebases for annotations. Existing DS-NER benchmark datasets use NER benchmark datasets to simulate the distant supervision setting by replacing the human annotations on training datasets with knowledgebase annotations Liang et al. (2020); Shang et al. (2018); Zhou et al. (2022). There are some potential biases of current DS-NER benchmarks. 1) only BC5CDR Shang et al. (2018) dataset comes from professional domains where DS-NER tasks are in high demand. 2) a series of hand-crafted procedures were applied to the current DS-NER benchmarks. Such procedures are entity-type dependent and require substantial human effort, and thus may not be generalizable to other DS-NER tasks. 3) The major entity types in existing DS-NER benchmarks consist of many proper nouns (such as Person’s names, location, and gene names), but in many applications, the entities to recognize need not to be proper nouns.
@@ 2.2 DS-NER Methods
Handling annotation errors in DS-NER tasks has drawn special attention. Here we briefly discuss a few representative approaches.
One line of work assumes that distant supervision often has high-quality positive labels, therefore focusing on alleviating the impact of false negative errors. Some methods address this issue by detecting potential entity candidates Shang et al. (2018); Xu et al. (2023). Some methods adopt positive and unlabeled learning to tackle false negative errors from the loss estimation perspective without separate detection steps Peng et al. (2019); Zhou et al. (2022). Due to its superiority in tolerating false negative errors, we embed Conf-MPU Zhou et al. (2022) into our proposed method.
Another line of work simultaneously considers annotation errors of all types. Some methods propose to train an initial model and apply a self-training framework to reduce the impact of noise Liang et al. (2020); Liu et al. (2021); Zhang et al. (2021b); Qu et al. (2023); Li et al. (2023). Methods such as RoSTER Meng et al. (2021) and SANTA Si et al. (2023) endeavor to diminish noise effects using loss functions tailored for noise resilience.
These DS-NER approaches obtain promising performance on existing DS-NER benchmark datasets. However, in a real-life DS-NER application, we observe that they fail to obtain satisfactory performances. These methods are trained on noisy labels initially, so the noise detection may have unknown biases and cause irreparable damage.
3 QTL Benchmark
We present QTL, a real-life DS-NER application in the animal science domain. The entity type to recognize is “trait”, an important task in the construction of genotype-phenotype databases for advancing livestock genomics research and breeding methodologies. Different from previous DS-NER datasets, where entities consist of many proper nouns, trait entities consist of descriptive expressions. For example, in the sentence “There are several reports on candidate gene analysis of subcutaneous fat depth traits using similar Wagyu × Limousin F2 populations”, “subcutaneous fat depth traits” is a trait entity that needs to be identified.
We collected a corpus with 1,716 abstracts from quantitative trait locus (QTL) studies related to six species, which were meticulously selected from PubMed [^1] https://pubmed.ncbi.nlm.nih.gov/ by domain experts. There are 18706 sentences and 514176 tokens in the train set. For the distant annotation process, we gathered a specialized dictionary of 3,884 trait names from four established domain ontologies [^2] Vertebrate Trait (VT) Ontology, Livestock Product Trait (LPT) Ontology, Livestock Breed Ontology (LBO), and Clinical Measurement Ontology (CMO). To evaluate DS-NER approaches, a well-trained domain curator provided ground truth annotations for 107 abstracts. These were then divided into validation and test sets, containing 21 and 1,044 sentences, 952 tokens and 32251 tokens, respectively. The test set has 1219 entities.
It is noteworthy that in the real-life application of QTL, the validation set used for hyperparameter tuning is quite small, comprising only 21 sentences. This limited size of the validation set may impact the tuning of hyperparameters during the model training process, potentially affecting the model’s performance. This issue reflects a realistic challenge encountered in DS-NER applications in real-life scenarios.
Metrics: In the QTL application, according to the curator’s practical needs, identifying potential entities is more important than identifying precise boundaries. Therefore, besides the strict span-level Precision (P), Recall (R), and F1 scores, we also introduce relaxed Precision (P), Recall (R), and F1 scores to evaluate the performance of DS-NER methods for practical usage. For relaxed metrics, it deems a predicted span correct if there is at least one overlapping word with the ground truth annotation.
@@ 4 Methodology
Figure 1: The CuPUL Framework
In DS-NER, we denote an input sentence with $M$ tokens as $\boldsymbol{x}=[x_{1},x_{2},\cdots,x_{M}]$ and denote corresponding annotations as $\boldsymbol{y}=[y_{1},y_{2},\cdots,y_{M}]$, $y_{i}\in\{0,1,\cdots,k\}$, where 0 denotes unlabeled types and $1,\cdots,k$ denote $k$ entity types. For the models, a pre-trained language model such as RoBERTa is used to encode token representations and it is followed by a softmax function to forward the prediction of entity labels for each token in the sentence.
Curriculum learning has two main steps: difficulty estimation and curriculum scheduler Kocmi and Bojar (2017). We have more discussion and related work in Appendix .1. Following the main steps of curriculum learning, we design the framework of CuPUL, which is shown in 1. The process starts by distantly labeling the corpus using knowledge bases, and then several voters are trained using this data to calculate token difficulty scores. Finally, CuPUL trains a NER classifier following the curriculum scheduler using confidence-based positive-unlabeled learning.
@@ 4.1 Difficulty Estimation
Motivated by the token-level noises in DS-NER tasks, we design the difficulty estimator and the curriculum scheduler at the token level as well. It allows the model to learn from one sentence by ignoring the noisy tokens. For example, in the sentence “Peter(PER) lives(O) in(O) America(ORG)”, “Peter”, “lives”, and “in” are clean samples, and “America” is a noise sample. The model can learn from “Peter lives in X” by ignoring the noise in the sentence. The token’s difficulty score reflects its inherent learnability. These scores are estimated using the disagreements between basic NER models or voters.
@@ 4.1.1 Voters
The design of the voters demands simplicity and variability. To balance efficiency and diversity, several voters are trained to increase prediction results variability. Label imbalance in NER tasks is mitigated by sampling from negative samples, introducing randomness to voters. Randomness is further enhanced with different random processes, such as shuffling and initialization.
For training the voters, a neural network for NER classification is used, as defined at the beginning of Section 4. The voters are trained using a regular multi-class classification risk function. The training process follows the Positive-Negative setting, where 0 represents non-entity type. This is doable because the performance of the voter itself does not affect the final outcome of CuPUL, which we will introduce in the section 4.2.
@@ 4.1.2 Difficulty Scores
Given that there are $V$ voters, the final difficulty score for each token $x$ is defined as the average of the non-identical pairs among all voters as:
@@ 4.2 Curriculum Design
To avoid overfitting of negative samples, we set the model learning process as Positive-Unlabeled (PU) learning, where data labeled with 0 is considered unlabeled rather than non-entity. PU learning assumes that the unlabeled data follows the distribution of the entire dataset Zhou et al. (2022). To fulfill the assumption, we directly incorporate all unlabeled data into the curriculum to prevent too less unlabeled tokens in curricula. Different curriculum partitions are executed solely on the labeled positive data. In this way, only the learning order of positive samples can affect CuPUL performance.
Our curriculum is built on token difficulty scores $H$, which are mostly "easy" as they follow a long-tail distribution (Figure 3). Previous work Platanios et al. (2019); Gnana Sheela and Deepa (2013) suggests that a curriculum with uniformly ranged difficulty scores might lead to most tokens belonging to the first curriculum, making curriculum learning ineffective. Hence, we propose using a power-law selector to construct a more effective curriculum scheduler.
To build the curricula, we first arrange $T_{u}$ unlabeled tokens followed by $T_{p}$ positive-labeled tokens sorted by their difficulty scores in ascending order. The first curriculum consists of all unlabeled tokens and the first $\tau T_{p}$ labeled positive tokens, where $\tau$ ($10<\tau<$ 1) is a selective factor. The second curriculum consists of the first $\tau^{2}T$ tokens from the remaining $(1-\tau)T_{p}$ tokens. This selection process continues until the penultimate curriculum. The remaining tokens are placed in the final curriculum. These curricula are denoted as $C_{1},C_{2},...,C_{\eta}$:
$\begin{split}C_{1}&:1\sim T_{u}+\tau T_{p}\\ C_{2}&:T_{u}+\tau T_{p}+1\sim T_{u}+(\tau+\tau^{2})T_{p}\\ &\cdots\\ C_{\eta}&:T_{u}+(\tau^{\eta-1}+\cdots+\tau)T_{p}+1\sim T_{u}+T_{p}.\\ \end{split}$
For example, suppose $T_{p}=20$, $T_{u}=80$, $\tau=0.5$, and $\eta=3$. Then, the first curriculum consists of tokens indexed from 1 to 90 ($80$ unlabeled tokens and the $10$ easiest positive tokens), the second curriculum consists of tokens indexed from 91 to 95, and the last curriculum consists of tokens indexed from 96 to 100.

### v3

@@ Abstract
Distantly-Supervised Named Entity Recognition (DS-NER) uses knowledge bases or dictionaries for annotations, reducing manual efforts but rely on large human labeled validation set. In this paper, we introduce a real-life DS-NER dataset, QTL, where the training data is annotated using domain dictionaries and the test data is annotated by domain experts. This dataset has a small validation set, reflecting real-life scenarios. Existing DS-NER approaches fail when applied to QTL, which motivate us to re-examine existing DS-NER approaches. We found that many of them rely on large validation sets and some used test set for tuning inappropriately. To solve this issue, we proposed a new approach, token-level Curriculum-based Positive-Unlabeled Learning (CuPUL), which uses curriculum learning to order training samples from easy to hard. This method stabilizes training, making it robust and effective on small validation sets. CuPUL also addresses false negative issues using the Positive-Unlabeled learning paradigm, demonstrating improved performance in real-life applications. [^1] Our code and new QTL dataset are available at https://github.com/liyp0095/CuPUL.
@@ A New Benchmark and a Simple Approach
Yuepei Li  Kang Zhou  Qiao Qiao  Qing Wang  Qi Li Iowa State University, Ames, Iowa, USA {liyp0095, kangzhou, qqiao1, qingwang, qli}@iastate.edu
@@ 1 Introduction
Distantly-Supervised Named Entity Recognition (DS-NER) is a task to leverage existing knowledge bases (KBs) or dictionaries to provide annotations for named entity recognition tasks. This approach significantly reduces the need for labor-intensive manual annotations, but it faces challenges due to issues in automated annotations, such as false positives and false negatives. To address the annotation errors, various methods are proposed. Some studies focus on false negative issues Shang et al. (2018); Peng et al. (2019); Zhou et al. (2022). Others propose to tackle general noisy annotations through noise removal processes Meng et al. (2021); Liang et al. (2020); Hedderich and Klakow (2018); Zhang et al. (2021a); Liu et al. (2021).
To assess the effectiveness of existing DS-NER approaches, we introduce a real-life DS-NER dataset, QTL (Quantitative Trait Locus), which is annotated for trait entities in the animal science domain. Unlike previous datasets, QTL contains a very small validation set of only 21 sentences, avoiding the significant manual effort required to obtain large validation sets in real-life scenarios. In contrast to previous benchmark datasets where entity mentions often comprise proper nouns, the trait entities in the QTL dataset are descriptive terms, such as “tail size” and “hoof color”.
While existing DS-NER methods perform well on benchmark datasets such as CoNLL2003, often rivaling fully supervised approaches, they consistently fail when applied to our QTL dataset. This motivates us to re-examine existing DS-NER approaches. We identify some issues: Some approaches Liang et al. (2020); Zhang et al. (2021b); Qu et al. (2023) deviate from the DS-NER framework and directly use the test set for hyperparameter tuning, leading to unreliable performance. Some approaches Shang et al. (2018); Meng et al. (2021) train their models with fixed hyperparameters, yet fail to achieve consistent results across different datasets. The remaining approaches Wang et al. (2023a); Wu et al. (2023) employ a validation set from fully supervised (FS) data for parameter tuning. These approaches overlook the significant manual labor required to obtain a validation set for parameter tuning in real-life scenarios, affecting the robustness of existing methodologies when applied to real-life applications with a small validation set, thereby compromising the reliability of these approaches.
@@
To solve the issues mentioned above, we present a simple yet effective approach inspired by Curriculum learning and Positive-Unlabeled (PU) learning , named CuPUL. The motivation behind curriculum learning is that deep learning models are non-convex and trained using batches of samples, so the order of training data can significantly impact model performance. Curriculum learning rearranges the batches of training samples such that the model learns from easy to hard samples and revisits easier samples more frequently. With this new arrangement, models tend to converge to a better local optimum. Furthermore, we design a token-level curriculum arrangement to address token-level noise in DS-NER tasks. We observe that "easy samples" are usually cleaner, and learning from these first can initially avoid label noise, making the model more robust. To tackle false negative issues, we adopt the Positive-Unlabeled learning paradigm.
@@
  • We present a real-life DS-NER dataset, QTL, and test the performance of the existing state-of-the-art methods. We observe that many methods do not follow the practical DS-NER setting and have unsatisfactory performance.
@@
  • We propose a simple method CuPUL to address the noise issue in DS-NER. We empirically demonstrate that CuPUL can significantly outperform the state-of-the-art DS-NER method on the QTL dataset and different benchmark datasets.
2 QTL Benchmark
@@
To reduce the cost of human-annotated training data for NER tasks, DS-NER uses professional dictionaries or knowledgebases for annotations. Existing DS-NER benchmark datasets use NER benchmark datasets to simulate the distant supervision setting by replacing the human annotations on training datasets with knowledge base annotations Liang et al. (2020); Shang et al. (2018); Zhou et al. (2022). Among these benchmarks, only the BC5CDR Shang et al. (2018); Li et al. (2016) dataset comes from professional domains where DS-NER tasks are in high demand.
We present Quantitative Trait Locus (QTL), a real-life DS-NER application in the animal science domain. The entity type to recognize is “trait”, an important task in the construction of genotype-phenotype databases for advancing livestock genomics research and breeding methodologies (examples in Table 1).
Compared to previous DS-NER benchmark datasets, the QTL dataset presents two main distinctions: 1) The entities in QTL are generally longer in length. 2) Unlike previous benchmarks, where entities are typically proper nouns, QTL entities are often descriptive expressions. To describe the distinct characteristics of trait entities, we compare the Trait type from the QTL test set with entity types (PER, LOC, ORG, MISC) from the CoNLL03 Tjong Kim Sang and De Meulder (2003) test set. Some key statistics are presented in Table 2, revealing that Trait entities have a longer entity length on average and a lower proper noun ratio compared to the benchmark dataset. These characteristics make the QTL dataset unique and introduce additional challenges in the DS-NER task.
No. Trait Entity Example
1 Fatty acid composition of milk
2 Dwarf phenotype
3 Total number of piglets born per litter
4 Body mass index
5 The number of first to third births
Table 1: List of Trait Entities from Test Set of QTL.
Dataset Entity Type # Entity Average Entity Length Proper Noun Ratio
PER 1617 1.71 0.97
CoNLL03 LOC 1662 1.16 0.98
ORG 1656 1.51 0.92
MISC 693 1.32 0.60
QTL Trait 1219 1.98 0.24
Table 2: Statistics on Entity Types of CoNLL2003 and QTL. The average entity length indicates the number of words per entity, and the proper noun ratio reflects the proportion of entities that contain at least one proper noun.
To establish the QTL dataset, we collected a corpus with 1,716 abstracts carefully selected from PubMed [^2] https://pubmed.ncbi.nlm.nih.gov/ by domain experts for QTL studies related to six species: cattle, pig, goat, sheep, chicken, and rainbow trout. We randomly selected 1,609 abstracts in this corpus to establish the training data. The training data consists of 18,706 sentences with 514,176 tokens. For the distant annotation process, the domain experts gathered a specialized dictionary of 3,884 trait names from four established domain ontologies [^3] Vertebrate Trait (VT) Ontology, Livestock Product Trait (LPT) Ontology, Livestock Breed Ontology (LBO), and Clinical Measurement Ontology (CMO). Examples can be found at https://www.animalgenome.org/QTLdb/export/trait_mappings. After obtaining the dictionary, string matching was used to distantly annotate the training corpus. Then from the 1609 papers, we randomly selected 21 sentences (with 952 tokens) and had a well-trained domain curator provide manual annotations to form a validation set.
This curator later provided annotations for the remaining 107 abstracts to form a test set, which covered all six species of interest. To assess annotation quality, we had a second domain curator check the annotations on 10 randomly selected abstracts. The two curators had a total agreement. Therefore, we used the annotations as ground truth. More annotation details are described in Appendix E. The test set contains 1,044 sentences with 32,251 tokens and 1,219 entities.
Notably, the validation set is quite small in the QTL dataset. This practice followed the motivation of DS-NER tasks, where the human effort should be minimized at the training stage. This limited size of the validation set may impact the tuning of hyperparameters during the model training process, potentially affecting the model’s performance. This issue reflects a realistic challenge encountered in DS-NER applications, which requires the model to be robust and not sensitive to hyperparameters.
Annotation Limitations: Due to the high cost, the majority of the annotations are provided by a single curator. One observation from the curators is that there is a considerable amount of discontinued trait entities. For example, in “milk protein, lactose, and fat percentage”, there are three entities: milk protein percentage, milk lactose percentage, and milk fat percentage. Due to the annotation software limitation, this example was annotated as “milk protein”, “lactose”, and “fat percentage”.
3 Related Work Analysis
Method Code Provided Code Runable Hyperparameter Tuning required Tuning Instruction Inference model Feasible
DS-NER without Self-training
AutoNER Fixed - Model at Final Epoch
Conf-MPU Not Fixed - Model at Final Epoch
MProto Not Fixed Model at Final Epoch
DS-NER with Self-training
BOND Not Fixed Best Model on Test
RoSTER Fixed - Model at Final Epoch
SCDL Not Fixed Best Model on Test
ATSEN Not Fixed Best Model on Test
DesERT Not Fixed First Student Model
Span-based DS-NER models
SANTA Not Fixed Model at Final Epoch
Top-Neg - - - -
CLIM - - - - -
Table 3: Feasibility Analysis of Exist Methods for DS-NER tasks.
@@ 3.1 DS-NER methods
We collected DS-NER methods published in major conferences in 2023 and their compared baselines. We categorize the existing DS-NER methods in three groups. 1) DS-NER with Self-training. To improve model performance, many DS-NER methods often incorporate a self-training step, utilizing mechanisms such as soft-label retraining and multi-model teacher-student frameworks. This group includes BOND Liang et al. (2020), RoSTER Meng et al. (2021), SCDL Zhang et al. (2021b), ATSEN Qu et al. (2023) and DesERT Wang et al. (2023a). 2) DS-NER without Self-training. This group of methods focuses on addressing the model’s effectiveness in handling noise or false negatives in DS-NER. While these methods can incorporate self-training mechanisms, it is not the primary focus of these methods. This group include AutoNER Shang et al. (2018), Conf-MPU Zhou et al. (2022), MProto Wu et al. (2023). 3) Span-based DS-NER. The final group of methods differs from the previous two, as it is based on span-based prediction rather than sequence labeling. These methods treat each span within a sentence as the prediction target. Previous work Li et al. (2023) has shown that span-based NER models often outperform sequence-based NER methods in terms of effectiveness, albeit at the cost of increased algorithmic complexity. This group includes Top-Neg Xu et al. (2023), CLIM Li et al. (2023) and SANTA Si et al. (2023). More details can be found in Appendix D.
3.2 Method Analysis
We first analyze the feasibility of existing DS-NER methods in real-life applications. For a method to be considered feasible, it must provide runnable code and instructions for hyperparameter tuning if necessary. Table 3 presents our feasibility analysis results base on the manuscripts and code repositories (accessed in April 2024). We find that 1) MProto and SANTA do not provide hyperparameter tuning instructions; 2) CLIM and Top-Neg do not provide runnable code; and 3) BOND, SCDL, and ATSEN selected their inference model based on performance on the test set according to their released repositories. Thus in our empirical studies, for a fair comparison, we only re-examine feasible methods and update some methods to select the inference model based on performance on the validation set only.
The motivation of DS-NER methods is that the manual annotations are too costly to obtain. Therefore, to reduce the amount of manual annotation, the annotations in the training set come from knowledge bases or dictionaries, and the manual labeled validation set should not be large either. Existing methods focus on the first setting while neglecting the importance of the second setting. We analyzed the feasible methods in Table 3 based on these DS-NER settings and have the following observations. First, AutoNER and RoSTER use fixed hyperparameters. These approaches do not require hyperparameter tuning, thereby avoiding the need for a validation set. Second, Conf-MPU provides a strategy for pre-selecting hyperparameters, so it does not require a validation set either. However, the remaining methods (BOND, SCDL, ATSEN, and DesSERT) need a validation set for hyperparameter tuning. The size of the validation set may affect their performance. We present a detailed analysis of this impact in Section 5.
We do not consider the few-shot NER setting in this paper. The recent SOTA performance of few-shot NER is achieved by LLMs with in-context learning (ICL), but it is unsatisfactory, particularly in specific professional domains Munnangi et al. (2024); Hu et al. (2024); Monajatipoor et al. (2024). Moreover, the high cost of LLMs makes it less feasible for large-scale annotation tasks. BERT-based DS-NER methods are easy to deploy and can achieve high performance with low cost. Therefore, we focus on the DS-NER setting using the BERT-based model.
@@ 4 Methodology
Figure 1: Overview of CuPUL
In this section, we introduce a simple DS-NER method that combines the advantages of curriculum learning and PU learning. Figure 1 shows the overview of the proposed method CuPUL. The method starts by training several voters using the distantly annotated data to calculate token difficulty scores. Then CuPUL trains a NER classifier following the curriculum scheduler using confidence-based positive-unlabeled learning risk estimation.
Problem Formulation: We denote an input sentence with $M$ tokens as $\boldsymbol{x}=[x_{1},x_{2},\cdots,x_{M}]$ and denote corresponding annotations as $\boldsymbol{y}=[y_{1},y_{2},\cdots,y_{M}]$, $y_{i}\in\{0,1,\cdots,k\}$, where 0 denotes the unlabeled type and $1,\cdots,k$ denote $k$ entity types. For the models, a pre-trained language model such as RoBERTa is used to encode token representations and followed by a softmax function to forward the prediction of entity labels for each token in the sentence.
@@ 4.1 Difficulty Estimation
Curriculum learning has two main steps: difficulty estimation and curriculum scheduler Kocmi and Bojar (2017). More details and related work of curriculum learning are discussed in Appendix A.
Motivated by the token-level noises in DS-NER tasks, we design the difficulty estimator and the curriculum scheduler at the token level as well. It allows the model to learn from one sentence by ignoring the noisy tokens. For example, in the sentence “Peter(PER) lives(O) in(O) America(ORG)”, “Peter”, “lives”, and “in” are clean samples, and “America” is a noisy sample. The model can learn from “Peter lives in X” by ignoring the noise in the sentence. The token’s difficulty score should reflect its inherent learnability. These scores are estimated using the disagreements between basic NER models or voters.
@@ 4.1.1 Voters
For training the voters, a neural network for NER classification is used. The design of the voters demands simplicity and variability. Thus, the voters are trained using a regular multi-class classification risk function. The training process follows the Positive-Negative setting, where 0 represents non-entity type. Label imbalance in NER tasks is mitigated by sampling negative samples. Note that the performance of the voter itself does not affect the final outcomes of CuPUL, which we will introduce in the section 4.2.
@@ 4.1.2 Difficulty Scores
Given that there are $V$ voters, the final difficulty score for each token $x$ is defined as the average of the non-identical pairs among all voters:
@@ 4.2 Curriculum Design
To avoid overfitting negative samples, we adopt Positive-Unlabeled (PU) learning based risk estimation, treating data labeled with 0 as unlabeled rather than non-entity. PU learning assumes the unlabeled data represents the entire dataset’s distribution Zhou et al. (2022, 2023). To meet this assumption, we include all unlabeled data in the first curriculum, scheduling only the labeled positive data.
Our curriculum is based on token difficulty scores $H$, which follow a long-tail distribution, making most tokens “easy” (Figure 2). Previous research Platanios et al. (2019); Gnana Sheela and Deepa (2013) indicates that a uniform difficulty range may render curriculum learning ineffective. Therefore, we propose a power-law selector for a more effective curriculum scheduler.
To build the curricula, we first arrange all $T_{u}$ unlabeled tokens followed by $T_{p}$ positive-labeled tokens sorted by their difficulty scores in ascending order. The first curriculum consists of all unlabeled tokens and the first $\tau T_{p}$ labeled positive tokens, where $\tau$ ($10<\tau<$ 1) is a selective factor. The second curriculum consists of the first $\tau^{2}T$ tokens from the remaining $(1-\tau)T_{p}$ tokens. This selection process continues until the penultimate curriculum. The remaining tokens are placed in the final curriculum. These curricula are denoted as $C_{1},C_{2},...,C_{\eta}$. For example, suppose $T_{p}=20$, $T_{u}=80$, $\tau=0.5$, and $\eta=3$. Then, $C_{1}$ consists of tokens indexed from 1 to 90 ($80$ unlabeled tokens and the $10$ easiest positive tokens), $C_{2}$ consists of tokens indexed from 91 to 95, and $C_{3}$ consists of tokens indexed from 96 to 100.